# Evolving Assembly Code in an Adversarial Environment

### v1

@@ 4.3. Genetic Operators
We used Koza’s standard mutation and crossover operators (Koza et al., 1994) that operate on the survivors’ parts, which are represented as trees. Specifically, we used the grow sub-tree (i.e., sub-part) mutation and the exchange sub-tree (sub-part) crossover. We added two more operators. The duplicate-tree (part) mutation takes the best tree (part) of a survivor and replaces the second part with it. The exchange-trees (parts) crossover replaces one of the trees (parts) of the first individual with one of the trees (parts) of the second.
5. Experiments and Results
@@
We carried out a comprehensive set of experiments aimed at winning the top human-written survivors. Our code is written in Python, using the EC-KitY toolkit (Sipper et al., 2023). Our code and data are at github.com/anonymous-submission. The code for the human-written survivors we compete with can be found at (Leshem and Eyzenberg, 2012).
@@
Experiments were conducted on a cluster of 96 nodes and 5,408 CPUs (the most powerful processors are AMD EPYC 7702P 64-core, although most have lesser specs). 64 CPUs and 150 GB RAM were allocated for each run to parallelize the evaluation. In practice, each run took approximately two days.
The specific hyperparameters utilized in the experiments and their chosen values are detailed in Table 1.
Table 1. Evolutionary hyper-parameters.
@@
Termination 2,000 generations or convergence with a winning strike
@@
The operators were sequentially applied to individuals with different probabilities (Table 1). The evolution was set to terminate when 2,000 generations are reached, or before, depending on whether convergence between best and average fitness values is achieved in addition to a monotonic non-increasing winning strike of 200 generations.
We repeated each experiment ten times to prove consistency. Our individuals’ average fitness and standard deviation against each of the past years’ winners are in Table 2. We consider an average engine score higher than 0.5 a winning result for our individual. Notably, evolution managed to evolve assembly programs, which won almost 78% of past years’ human-written winners.
In the following sections, we inspect the code of the evolved solvers. The inspection revealed that the evolved survivors managed to win complex and long survivors using a relatively small code fragment. Although GP frequently evolves long code, sometimes only a small part of it is used in the program run flow and yet manages to win. As we will demonstrate, this shows how evolution found the Achilles’ heel in the opponents’ code and utilized it for its benefit.
Table 2. Test average fitness and standard deviation over ten experiments of our best individuals against past years’ winners.
@@
Figure 4. Evolved survivor against Zorg (2012 winner). The two parts of the evolved survivor utilized Zorg’s Achilles’ heel by writing data to a part of its program. Strike-through text denotes run-time changed code.
5.1. Qualitative Analysis of Evolved Survivors
5.1.1. Utilizing Achilles’ heel
@@
One of the clearest examples is Zorg—the 2012 winner. Zorg writes an important code fragment for its future run on memory address zero. The evolution process noticed it in about 100 generations and overridden this memory by addressing di, which holds the value zero, depriving Zorg from winning (see line 14 in both parts of Figure 4, which includes the dw translation to assembly commands and the effect on the following commands). Zorg’s code is significantly longer and more complicated than the evolved fragment that overtook it. The evolved survivor manages to win Zorg in about 70% of the cases despite the weakness finding due to the randomness in the game’s execution order.
5.1.2. Concentrated vs. Scattered Memory Writes
@@
During evolution, we noticed several spikes in the best fitness. For example, when training against BlocksOfGuru, there were spikes in the fitness of the best individuals in generations 206 and 256 (see Figure 5). To analyze these spikes, we ran a game with BlocksOfGuru against these individuals and the overall best individual (from generation 1,769). The results and memory image are depicted in Figure 6. We can see that most memory writes were made by the second part of the 1,769 and 256 individuals that cover the arena with scattered green and pink dots. 206’s second part performed less, yet a significant number of writes in yellow are concentrated in several areas. All of the first part performed little to no new memory writes. Inspecting their cleared runnable code (Listings 12 and 15) reveals that the first parts of 256 and 1,769 run in the loop written in their bottom, which keeps the individual alive but does not perform attacking actions—as seen in the lack of their color in the arena. In 206, both parts run in a loop due to jmp ax at the end, which jumps into the beginning of the code. Most memory writes of all individuals are performed using addressing si, yet with adding different constants to si. 256 and 1,769 use special constants like 65,535 and @start, while 206 does not. The first two exceed the bounds of word data, and the computation is thus written to an address defined as the special constant modulo $2^{16}$, which results in scattered writes. The evolutionary process discovered that scattered writing has more chances to encounter adversary code, as reflected in their higher scores and it is reflected in runs against other survivors as well.
Figure 5. BlocksOfGuru 2021
@@
Figure 6. The memory image of BlocksOfGuru vs. best individuals from generations 206, 256, and 1,769. The scattered green and pink dots are memory bytes written by the 1,769 and 256 individuals, respectively. Non-reachable code is stroke through.
(a) Evolution vs. GreeniEs (2020)
(b) Evolution vs. Zeus (2006)
Figure 7. Vertical vs. horizontal memory write. Horizontal writings of evolved survivors (in purple and yellow) “cut” the human-written adversaries that write in horizontal lines (green and red) by writing on their code before they reach their code.
@@
Figure 11. 256 part 2
Figure 12. Comparing the code of best individuals against the BlocksOfGuru survivor.
@@
Figure 14. 1,769 part 2
Figure 15. Comparing the code of best individuals against the BlocksOfGuru survivor.
5.1.3. Vertical vs. Horizontal Memory Writes
Another interesting pattern we detected in evolved survivors was writing vertical bytes into memory. That is, sequentially writing a byte every 256 bytes, creating a vertical line in the memory state. This contrasts with human-written survivors, who usually write memory in horizontal lines (consecutive bytes). As a result, the evolved survivors were able to “cut” the adversary by writing on its code before the adversary reached its code. This is depicted, for example, in the memory state of our evolved individuals against Zeus (2006) and GreeniEs (2020) (see Figure 7). The evolved survivors, in purple and yellow, write their code in vertical lines, thus cutting their adversaries’ horizontal writing (in red and green). Writing vertically assures reaching the opponent’s code faster because most submitted survivors take advantage of the maximum allowed code size, thus filling at least one memory row entirely. Therefore, filling one or a few columns will be faster than filling complete rows. A similar pattern was found in a significant part of runs against other survivors.
@@
5.1.4. Random Generator Pattern
The original program had difficulties overtaking a few previous years’ winners, specifically FSM 2010, IamAA 2014, LoudBugFix 2016, and TheHeapMen 2022, resulting in an average score lower than 0.5. We assumed that human-written survivors may use randomness to be unpredictable, while our BNF has no element of randomness. Thus, we decided to add Pseudo-Random Number Generator (PRNG) patterns to our BNF. We added Linear Congruentional Generator (LCG) and XOR-Shift Generators implementation to our grammar as shown in LABEL:table3_random_set and ran the evolution against the above adversaries again for ten runs each.
As shown in Table 3, using randomness improved the number of games evolution won and the average score in 3/4 of the cases. The majority of the best evolved individuals contained at least one of the random patterns. However, in some, the pattern appears in an unreachable code segment or outside the loop, meaning it only executes once. We believe it helped the evolution process, even though the winning survivor does not actively use it. The use of randomness enhanced the use of scattered writing patterns (see Section 5.1.2) for some of the survivors and evolved a combined horizontal-vertical writing pattern for others, in contrast to the described in Section 5.1.3. The innovation the random pattern caused is the combination of the described patterns together, resulting in scattered writing in horizontal lines that expand vertically, as seen in Figure 16. We can see the yellow and purple memory cells that are being filled horizontally at the beginning. Afterwards, the created lines expand vertically, and everything is done using scattered writing.
@@ 5.1.4. Random Generator Pattern
Table 3. Test game result with and without randomness.
@@
Figure 16. Scattered, horizontal and vertical writing in the survivor which evolved using randomness patterns against FSM (2010).
5.2. Comparison to LLMs
We compared our results to ChatGPT, one of today’s leading LLMs. We explained the CodeGuru game, the special opcodes, and the rules using prompt engineering. We asked the chat for a survivor with the best chance to overtake others. The suggested code was tested against simple competition survivors that are given as examples in the game’s web engine. At first, the provided survivors did not compile. After supplying corrections, they compiled; however, they achieved poor scores. Even when we supplied the opponents’ code to ChatGPT, its produced survivor did not win. The prompts we used can be found at https://chat.openai.com/share/5ebda97c-06f8-4430-bd04-78a5abfc74ea.
6. Conclusion
This work presented the evolution of assembly code from scratch in an adversarial environment, specifically the CodeGuru Xtreme competition. Through the use of GP, we synthesized assembly code that outperformed human-written winning survivors from past years. Our evolved programs were able to identify and exploit weaknesses in the programs they were trained against. Additionally, we compared our approach with the use of a Large-Language Model (LLM) and demonstrated that the LLM was unable to generate a survivor capable of winning any competition.
The CodeGuru Xtreme competition serves as a valuable platform for studying GP and code evolution in adversarial environments. To facilitate further research in this direction, we have provided a comprehensive qualitative analysis of the evolved survivors and the weaknesses they have identified. While we were able to overtake most of the top human-written survivors, the domain is far from being solved. There are still some survivors that evolution failed to win, and the evolved survivors were only able to win against one survivor at a time.
The implications of this work extend beyond the CodeGuru Xtreme competition. Utilizing evolution to detect weaknesses in other survivors has important applications in cyber-security. Furthermore, the assembly language used in this research is also relevant in the context of computer viruses. We believe that by modifying the fitness function, our approach can be employed to identify weaknesses in code and aid in their resolution.

### v2

@@ 4.3 Genetic Operators
We used Koza’s standard mutation and crossover operators koza1994genetic that operate on the survivors’ parts, which are represented as trees. Specifically, we used the grow sub-tree (i.e., sub-part) mutation and the exchange sub-tree (sub-part) crossover. We added two more operators. The duplicate-tree (part) mutation takes the best tree (part) of a survivor and replaces the second part with it. The exchange-trees (parts) crossover replaces one of the trees (parts) of the first individual with one of the trees (parts) of the second.
4.4 Improvements
Below are several improvements to the basic setup that we tested.
4.4.1 Random Generator Pattern
We wish to add randomness to the BNF to allow our survivors to be unpredictable. Thus, we add Pseudo-Random Number Generator (PRNG) patterns to our BNF. Specifically, we added Linear Congruential Generator (LCG) and XOR-Shift Generators implementation to our grammar as shown in LABEL:table3_random_set in the appendix.
4.4.2 Fitness Approximation
One of our main hold-backs is the fitness calculation time. Raising the Java engine, running 200 battles, and outputting the results into a file takes non-negligible time, which prolongs the evolution that itself requires many hours to complete a run due to G3P’s stochastic nature. To handle this issue, we combined a machine learning model, as presented in Tzruia2023Fitness, that learns how to approximate the fitness value without evaluating the individuals. Since their approach is designed for vector-based representation, we first create a mapping between our AST representation into a float vector form. For that, we use the AST’s size and the engine’s parameters (elaborated in Section 4.2). For example, if an individual has reached a score of 0.3, a lifetime of 3.21827, wrote 0.69897 bytes in a rate of 0.01, and its larger tree contains 25 nodes, its vector representation will be [0.3, 3.21827, 0.69897, 0.01, 25]. During evolution, the individuals’ vector representations and their actual fitness value are collected and used for training the ML model. The ML model switches between actual and approximated fitness according to defined conditions. In the actual fitness phase, it performs learning on truly evaluated individuals until it reaches a sufficient level of correct prediction and switches to approximation. In our case, we chose the switch condition to be based on cross-validation (CV) error as it reaches 5% of the maximal possible fitness, which equals 0.125 out of 2.5.
In each generation in the approximation phase, a certain percentage of the population is still evaluated regularly to maintain the model and save the accurate results of the best individual evolved so far. The percent which performed the best was 30%. Less sampling resulted in non-convergence of the evolution, and more sampling increased the evolution time.
4.4.3 Memetic Operators
As we will show, the use of fitness approximation dramatically reduced the computation time. Thus, we decided to utilize the fitness approximation to create smart memetic operators that use local search for further improving evolution.
Whenever we apply the basic genetic operators described above, we run them five times and choose the best individual according to the approximated fitness.
5 Experiments and Results
@@
We carried out a comprehensive set of experiments aimed at winning the top human-written survivors. Our code is written in Python, using the EC-KitY toolkit Sipper2023ECKity. Our code and data are at Assembly code generation. The code for the human-written survivors we compete with can be found at codeguru_repo.
@@
Experiments were conducted on a shared cluster of 96 nodes and a total of 5,408 CPUs (the most powerful processors are AMD EPYC 7702P 64-core, although most have lesser specs). 64 CPUs and 150 GB RAM were allocated for each evolutionary run (against one human adversary) to parallelize the evaluation. In practice, each run without the fitness approximation took approximately two days.
The specific hyperparameters utilized in the experiments and their chosen values are detailed in Table 1. The population size was chosen to be 192 to optimize the need for diversity, considering the resources of 64 CPUs and assigned time per evolutionary run. The Grow mutation probability was chosen to be 0.7 as the experiments that were conducted showed a clear tendency to better results with a high mutation rate, yet this rate allows evolution to perform a significant learning process.
Table 1: Evolutionary hyper-parameters.
@@
Termination 2,000 generations or convergence
with a winning strike
@@
Table 2: Test average fitness and standard deviation over ten experiments of our best individuals against past years’ winners.
@@
The operators were sequentially applied to individuals with different probabilities (Table 1). The evolution was set to terminate when 2K generations were reached, or before, depending on whether convergence between best and average fitness values was achieved in addition to a monotonic non-increasing winning strike of 200 generations.
We repeated each experiment ten times to test consistency. Our individuals’ average fitness and standard deviation against each of the past years’ winners are in Table 2. We consider an average engine score higher than 0.5 a winning result for our individual. Notably, evolution managed to evolve Assembly programs, which won almost 78% of past years’ human-written winners.
@@
Figure 4: Evolved survivor against Zorg (2012 winner). The two parts of the evolved survivor utilized Zorg’s Achilles’ heel by writing data to a part of its program. Strike-through text denotes run-time changed code.
5.1 Qualitative Analysis of Evolved Survivors
In this section, we inspect the code of the evolved solvers. The inspection reveals that the evolved survivors managed to win complex and long survivors using a relatively small code fragment. Although GP frequently evolves long code, sometimes only a small part of it is used in the program run flow and yet manages to win. As we will demonstrate, this shows how evolution found the Achilles’ heel in the opponents’ code and utilized it for its benefit.
5.1.1 Utilizing Achilles’ heel
@@
One of the clearest examples is Zorg—the 2012 winner. Zorg writes an important code fragment for its future run on memory address zero. The evolution process noticed it in about 100 generations and overridden this memory by addressing di, which holds the value zero, depriving Zorg of winning (see line 14 in both parts of Figure 4, which includes the dw translation to Assembly commands and the effect on the following commands). Zorg’s code is significantly longer and more complicated than the evolved fragment that overtook it. The evolved survivor manages to win Zorg in about 70% of the battles in a game (according to the average engine’s score) despite the weakness finding due to the randomness in the game’s execution order.
5.1.2 Concentrated vs. Scattered Memory Writes
@@
During evolution, we noticed several spikes in the best fitness. For example, when training against BlocksOfGuru, there were spikes in the fitness of the best individuals in generations 206 and 256 (see Figure 5). To analyze these spikes, we ran a game with BlocksOfGuru against these individuals and the overall best individual (from generation 1,769). The results and memory image are depicted in Figure 6. We can see that most memory writes were made by the second part of the 1,769 and 256 individuals that cover the arena with scattered green and pink dots. 206’s second part performed less, yet a significant number of writes in yellow are concentrated in several areas. All of the first part performed little to no new memory writes. Inspecting their cleared runnable code (Listings 11 and 14) reveals that the first parts of 256 and 1,769 run in the loop written in their bottom, which keeps the individual alive but does not perform attacking actions—as seen in the lack of their color in the arena. In 206, both parts run in a loop due to jmp ax at the end, which jumps into the beginning of the code. Most memory writes of all individuals are performed using addressing si, yet with adding different constants to si. 256 and 1,769 use special constants like 65,535 and @start, while 206 does not. The first two exceed the bounds of word data, and the computation is thus written to an address defined as the special constant modulo $2^{16}$, which results in scattered writes. The evolutionary process discovered that scattered writing has more chances to encounter adversary code, as reflected in their higher scores, and it is reflected in runs against other survivors as well.
Figure 5: BlocksOfGuru (2021)
@@
Figure 6: The memory image of BlocksOfGuru vs. best individuals from generations 206, 256, and 1,769. The scattered green and pink dots are memory bytes written by the 1,769 and 256 individuals, respectively. The concentrated yellow dots are memory bytes written by the 206 individual.
@@
Figure 10: 256 part 2
Figure 11: Comparing the code of best individuals against the BlocksOfGuru survivor.
@@
Figure 13: 1,769 part 2
Figure 14: Comparing the code of best individuals against the BlocksOfGuru survivor.
(a) Evolution vs. GreeniEs (2020)
(b) Evolution vs. Zeus (2006)
Figure 15: Vertical vs. horizontal memory write. Horizontal writings of evolved survivors (in purple and yellow) “cut” the human-written adversaries that write in horizontal lines (green and red) by writing on their code before they reach their code.
5.1.3 Vertical vs. Horizontal Memory Writes
Another interesting pattern we detected in evolved survivors was writing vertical bytes into memory. That is, sequentially writing a byte every 256 bytes, creating a vertical line in the memory state. This contrasts with human-written survivors, who usually write memory in horizontal lines (consecutive bytes). As a result, the evolved survivors were able to “cut” the adversary by writing on its code before the adversary reached its code. This is depicted, for example, in the memory state of our evolved individuals against Zeus (2006) and GreeniEs (2020) (see Figure 15). The evolved survivors, in purple and yellow, write their code in vertical lines, thus cutting their adversaries’ horizontal writing (in red and green). Writing vertically assures reaching the opponent’s code faster because most submitted survivors take advantage of the maximum allowed code size, thus filling at least one memory row entirely. Therefore, filling one or a few columns will be faster than filling complete rows. A similar pattern was found in many runs against other survivors.
@@
5.2 Random Generator Pattern
The original program had difficulties overtaking a few previous years’ winners, specifically FSM 2010, IamAA 2014, LoudBugFix 2016, and TheHeapMen 2022, resulting in an average score lower than 0.5. We assumed that the BNF extension of random generator patterns may improve our evolved survivors. We ran the evolution against the above adversaries again for ten runs each.
As Table 3 shows, using randomness improved the number of games evolution won and the average score in three out of four cases. The majority of the best-evolved individuals contained at least one of the random patterns. However, in some, the pattern appears in an unreachable code segment or outside the loop, meaning it only executes once. We believe it helped the evolution process, even though the winning survivor does not actively use it.
The use of randomness enhanced the use of scattered writing patterns for some of the survivors and evolved a combined horizontal-vertical writing pattern for others, in contrast to the vertical-only pattern. The random pattern allowed for the combination of the described patterns together, resulting in scattered writing in horizontal lines that expand vertically, as seen in Figure 16. We can see the yellow and purple memory cells that are being filled horizontally at the beginning. Afterward, the created lines expand vertically, and everything is done using scattered writing.
@@ 5.2 Random Generator Pattern
Table 3: Test game results with and without randomness.
@@
Figure 16: Scattered, horizontal, and vertical writing in the survivor, which evolved using randomness patterns against FSM (2010).
5.3 Fitness Approximation and Memetic Operators
For the ML-based fitness approximation approach (Section 4.4.2), we used the Ridge model over Lasso (see Tzruia2023Fitness) as it helps reduce overfitting that results from model complexity and doesn’t set the value of the coefficient to absolute zero. The complete fitness-approximation hyperparameters are given in Table 4, and they are used with the parameters described in Table 1.
The combination of the approximation model in our evolution led to about a 30% decrease in total evolution time while preserving the winning results achieved before (see Table 2). For example, the evolution vs Mamaliga (2011) previously took about 77 hours and was decreased to 56 hours with the same amount of generation until winning was achieved.
Table 4: ML hyper-parameters.
switch_condition CV Error
switch_threshold 0.125
sample_rate 30%
model Ridge
alpha 0.3
gen_weight sqrt
As the approximated fitness was relatively accurate, we decided to utilize it against the human-written survivors we were not able to overtake Table 3. We did that using the memetic operators described in Section 4.4.3, running 10 evolutionary processes against each opponent to establish our results, similar to the method described in Section 4.2. We also used the random-pattern extension in these experiments.
As Table 5 shows, there was a great improvement, and a win was achieved against all human-written survivors we were not able to win before. Furthermore, in all games in the test phase that resulted in winning, they achieved the best possible score of 1.0. They did not win all games, hence the average score of 0.8 and 0.9.
Table 5: Test game results with the memetic operators and random patterns.
Year Human Survivor w/o memetic operators w/ memetic operators
Avg. Score #Victories Avg. Score #Victories
2010 FSM 0.481 3/10 0.8 8/10
2014 IamAA 0.478 6/10 0.9 9/10
2016 LoudBugFix 0.402 2/10 0.9 9/10
2022 TheHeapMen 0.494 4/10 1 10/10
During the experiment, the ML model was able to learn the parameter weights in each run. The highest weight was given consistently to the score parameter, as expected, and the size parameter was approximately ignored in all cases. We also noticed that the writing_rate weight varied among the survivors, while the other weights had close values.
6 Application to Cyber Security
Signatures are widely used in anti-virus programs as a technique to detect known malware. They work by comparing specific patterns or characteristics of files against a database of known malware signatures. If a match is found, the anti-virus program can take appropriate action to quarantine or remove the malicious file. To bypass those techniques, malware developers have devised methods to alter their code in a manner that avoids detection by the signatures while still maintaining the malware’s functionality. This perpetuates the ongoing battle between protectors and attackers in an endless cycle.
The adversarial environment of CodeGuru is an interesting platform for analyzing the ability of evolution to overcome tools designed especially against it, like the ability of malware developers to overcome signatures. As mentioned in Section 3, there is a special opcode INT 0x87 in the game. It looks for a 4-byte sequence identical to the values stored in the registers AX:DX and replaces them with the values stored in BX:CX. The search is performed in the memory starting from address DI:ES moving up or down, determined by the direction flag. Therefore, if a command from a survivor is stored in AX:DX, it can be found and replaced by the adversary to an illegal command stored in BX:CX, causing the survivor to run it and be disqualified. This simulates the identification of malware by a signature.
We conducted an experiment utilizing the INT 0x87 command to simulate signature-based anti-virus. The adversaries chosen were XLII (2009) and GreeniEs (2020), both of which we had previously defeated and that incorporate INT 0x87 in their code. We subjected them to a complete evolutionary process, with a minor adjustment of halting evolution once the evolved population achieved a consistent win, indicated by an average fitness of 1.1, signifying a winning score (refer to Section 4.2). During this halt, we manually altered the adversaries’ AX:DX values to match the commands relied upon by the best-evolved individuals. Subsequently, we resumed the evolution, introducing the evolved population to a bespoke adversary.
In both cases shown in Figure 17, the fitness initially dropped in the following generation due to encountering an improved adversary. However, as evolution progressed, it managed to recover and even improve the fitness results to levels achieved before. The evolutionary process successfully replaced the targeted commands with different ones that exhibited similar behavior. These results demonstrate the capability of evolution to confront tools specifically designed to counter it. Furthermore, the use of a domain-independent Assembly grammar, coupled with the prevalence of malware written in Assembly, suggests that these findings could be leveraged to modify malware for evading signature-based anti-virus systems.
(a) Evolution vs. modified XLII (2009)
(b) Evolution vs. modified GreeniEs (2020)
Figure 17: The fitness dropped in generation 241 and 38, respectively, after the tailor-made adversary was created.
7 Conclusion
This work focused on evolving Assembly code for the CodeGuru competition with the objective of creating a survivor program that can run the longest in shared memory while withstanding attacks from adversary survivors. By defining a Backus Normal Form (BNF) for the Assembly language and employing Genetic Programming (GP) to synthesize the code, the study aimed to develop top-notch solvers. The evaluation of the evolved programs involved running CodeGuru games against human-written winning survivors, leading to the identification and exploitation of weaknesses in the opponent programs.
To enhance the evolutionary process, we presented three enhancements. Specifically, we added random patterns to the BNF to allow our survivors to be less predictable; we dramatically cut the overall training time approximating the fitness using a machine-learning (ML) approach; and finally, by developing ML-based memetic operators that allowed our survivors to outperform all pervious-years human-written survivors.
The significance of this research extends to the realm of cyber-security, where the evolved Assembly programs were adept at detecting vulnerabilities in the opponent survivors, showcasing the potential for utilizing evolutionary algorithms to identify and rectify code weaknesses. Our domain-independence Assembly BNF opens up possibilities for adapting the approach to various contexts by adjusting the fitness function to target specific code vulnerabilities.
Moreover, the CodeGuru competition serves as a valuable platform for studying Genetic Programming and code evolution within adversarial environments. The thorough qualitative analysis conducted on the evolved survivors and the vulnerabilities uncovered contributes to the body of knowledge in this area. By shedding light on the efficacy of evolutionary techniques in enhancing program robustness and resilience against attacks, this research paves the way for further investigations into evolutionary strategies for cybersecurity applications and code optimization in competitive settings.